# UAVDB: Trajectory-Guided Adaptable Bounding Boxes for UAV Detection

### v4

@@ 1 Introduction
Precise UAV detection is critical for effective monitoring and threat response. While modern object detection algorithms, such as YOLO-series detectors [11, 23, 22, 10] and transformer-based models [31, 2], have significantly advanced UAV detection, their performance is highly dependent on high-quality annotations. Without accurate, well-annotated datasets, even SOTA models struggle with real-world UAV detection, particularly for tiny UAVs. Existing UAV datasets can be broadly categorized into two types. The first type is ground-target UAV datasets, where UAV-mounted cameras capture objects like vehicles or pedestrians on the ground [24, 25, 7, 30, 18] and the second type is UAV-target datasets, where the UAV itself is the detection target. The latter can be further divided into three categories: 1) RGB frame with fixed-camera setup, where the camera remains stationary as presented in [21, 17], 2) RGB image with moving-camera setup, where the camera equipped on the UAV such as [19, 14], and 3) Infrared image UAV datasets, including single-frame datasets [6, 4, 5], and video-based Anti-UAV datasets [8, 9, 28, 29] which has been featured in four major challenge events.
@@ 2. $2$
    Provide the experiments validating PIC’s efficiency in terms of IoU and runtime, along with a comprehensive benchmark of UAVDB using SOTA YOLO-series detectors, including YOLOv8 [11], YOLOv9 [23], YOLOv10 [22], and YOLO11 [10].
@@ 3.1.2 Iterative Expansion
The bounding box expands outward by a fixed step $\delta$ in all directions per iteration:
$w_{t+1}=w_{t}+\delta,\quad h_{t+1}=h_{t}+\delta.$
@@ 3.1.3 Patch Intensity Calculation
At each step, the mean pixel intensity inside the bounding box is computed as:
@@ 4.1 Annotation Accuracy and Runtime Efficiency
First, human-labeled bounding boxes are considered ground truth annotations. For the fixed-size and thresholding [1] approaches, we use a 50$\times$50 region size and set the threshold value to 150, based on empirical testing, which yields the best performance. For the GrabCut [20] and Segment Anything Model (SAM) [13] approaches, we use the OpenCV package and the ViT-B pre-trained SAM model, respectively. As shown in Tab. 3, the PIC approach outperforms other methods, achieving the highest IoU while maintaining a minimal runtime of 0.007 seconds, comparable to the fixed-size method. This demonstrates that the computational time of the PIC process is negligible compared to image reading and output. In contrast, despite its advanced segmentation capabilities, SAM struggles with UAV-specific challenges, resulting in the lowest IoU. This illustrates that SAM cannot generalize effectively without retraining on a specific dataset. These results highlight the effectiveness of PIC in providing both accurate and computationally efficient UAV bounding box extraction, making it ideal for large-scale and real-time applications.
@@ 4.2 Benchmark on UAVDB
To benchmark the proposed UAVDB, we evaluate several SOTA detectors, including YOLOv8 [11], YOLOv9 [23], YOLOv10 [22], and YOLO11 [10]. These detectors were trained on 10,763 training images, 2,720 validation images, and 4,578 test images from UAVDB, as shown in Tab. 2. The experiments were conducted on a high-performance computing (HPC) system [16], utilizing an NVIDIA A100 GPU with 80 GB of memory. All models were trained with an image size of 640, a batch size of 32, 100 epochs, and eight workers. Mosaic augmentation was applied throughout training, excluding the final 10 epochs. Additionally, we fine-tuned the models using officially released pre-trained weights. Tab. 4 summarizes the training time, inference time, number of model parameters, FLOPs, and average precision (AP) for both validation and test results for each detector. Further, validation performance across epochs is illustrated in Fig. 4.
@@
Model Training Time Inference Time #Param. FLOPs $\text{AP}^{val}_{50}$ $\text{AP}^{val}_{50-95}$ $\text{AP}^{test}_{50}$ $\text{AP}^{test}_{50-95}$ (hours:mins:sec) (per image, ms) (M) (G) YOLOv8n 01:40:31 0.9 2.7 6.8 0.829 0.522 0.789 0.450 YOLOv8s 01:55:05 1.2 9.8 23.3 0.814 0.545 0.796 0.450 YOLOv8m 02:43:08 1.8 23.2 67.4 0.809 0.538 0.827 0.526 YOLOv8l 03:54:44 2.6 39.4 145.2 0.830 0.563 0.836 0.544 YOLOv8x 04:33:08 3.5 61.6 226.7 0.820 0.554 0.728 0.448 YOLOv9t 02:53:11 2.5 2.6 10.7 0.839 0.501 0.848 0.508 YOLOv9s 03:05:02 2.6 9.6 38.7 0.819 0.517 0.834 0.484 YOLOv9m 05:08:28 4.1 32.6 130.7 0.840 0.507 0.858 0.522 YOLOv9c 06:17:08 5.3 50.7 236.6 0.851 0.544 0.851 0.504 YOLOv9e 08:00:05 6.6 68.5 240.7 0.755 0.414 0.768 0.383 YOLOv10n 02:05:39 0.7 2.7 8.2 0.764 0.492 0.731 0.417 YOLOv10s 02:23:03 1.2 8.0 24.4 0.817 0.530 0.823 0.516 YOLOv10m 03:06:59 1.8 16.5 63.4 0.798 0.531 0.821 0.536 YOLOv10b 03:29:18 2.1 20.4 97.9 0.801 0.517 0.760 0.467 YOLOv10l 04:04:22 2.5 25.7 126.3 0.774 0.502 0.842 0.517 YOLOv10x 05:14:07 3.5 31.6 169.8 0.771 0.507 0.693 0.431 YOLO11n 01:50:00 0.9 2.6 6.3 0.847 0.527 0.856 0.539 YOLO11s 02:07:01 1.2 9.4 21.3 0.826 0.553 0.885 0.578 YOLO11m 03:07:40 1.9 20.0 67.6 0.827 0.588 0.843 0.578 YOLO11l 04:09:45 2.4 25.3 86.6 0.810 0.555 0.798 0.517 YOLO11x 05:20:38 3.6 56.8 194.4 0.812 0.560 0.782 0.534
Table 4: Performance metrics of YOLOv8 [11], YOLOv9 [23], YOLOv10 [22], and YOLO11 [10] models trained on UAVDB.
@@
Fig. 5 showcases the predictions made by the YOLO11s model on Dataset 5, which includes scenarios not present in the training data, highlighting the model’s ability to handle previously unseen situations. The detection results closely match the UAV sizes, validating the high fidelity of the bounding box annotations in UAVDB. Incorporating these high-quality predicted bounding boxes into the training set can further enhance the model’s capability to detect UAVs in various scenarios.
Figure 4: Validation performance of YOLOv8 [11], YOLOv9 [23], YOLOv10 [22], and YOLO11 [10] models over training epochs.
@@ 4.3 Discussion
As shown in Tab. 3, the human-labeled approach requires an average of 19 seconds per annotation, including zooming and detailed adjustments, which is time-consuming and labor-intensive, particularly for large datasets with tiny objects. In contrast, our PIC approach is approximately 2700$\times$ faster, though with a slight sacrifice in accuracy. Despite this, the proposed dataset has proven effective in training detectors, as demonstrated by the visualizations in Fig. 5. While the PIC method performs well on current datasets, we acknowledge that low-altitude UAV flights, with cluttered and rapidly changing backgrounds, may pose challenges. In such dynamic environments, local intensity variations could impact the pixel intensity metric for bounding box expansion. However, the adaptive nature of PIC, focusing on local intensity changes, allows it to handle moderate background texture variations. Further improvements could be explored for highly dynamic scenarios, such as incorporating multi-frame temporal information or background subtraction, to enhance robustness further and ensure consistent performance for the PIC technique.
@@ 5 Conclusion
In this paper, we introduced UAVDB, a high-quality dataset with precise bounding box annotations facilitated by our proposed PIC technique. PIC is an intuitive, efficient, and innovative method that enables automated, high-accuracy spatial annotations without manual labeling effort. UAVDB addresses critical limitations in existing UAV datasets, such as the lack of accurate annotations and insufficient environmental diversity, significantly enhancing the applicability of detection algorithms in real-world scenarios. Through IoU and runtime evaluations, alongside benchmarking SOTA YOLO detectors, we demonstrated the versatility of both UAVDB and the PIC approach under varied conditions. These results validate UAVDB as a valuable resource for advancing UAV detection technologies. Looking forward, the adaptability of PIC presents exciting opportunities for future work. Its lightweight design could be further enhanced by incorporating multi-frame temporal information or background subtraction to improve robustness in dynamic environments. The flexibility to fine-tune PIC for specific domains ensures its continued scalability and relevance across various UAV detection applications. As UAV detection technology progresses, UAVDB and PIC provide a solid foundation for advancing real-time, large-scale UAV detection in diverse environments.

### v5

@@ 1 Introduction
Precise UAV detection is critical for effective monitoring and threat response. While modern object detection algorithms, such as YOLO-series detectors [11, 24, 23, 10, 22] and transformer-based models [32, 2], have significantly advanced UAV detection, their performance is highly dependent on high-quality annotations. Without accurate, well-annotated datasets, even SOTA models struggle with real-world UAV detection, particularly for tiny UAVs. Existing UAV datasets can be broadly categorized into two types. The first type is ground-target UAV datasets, where UAV-mounted cameras capture objects like vehicles or pedestrians on the ground [25, 26, 7, 31, 18] and the second type is UAV-target datasets, where the UAV itself is the detection target. The latter can be further divided into three categories: 1) RGB frame with fixed-camera setup, where the camera remains stationary as presented in [21, 17], 2) RGB image with moving-camera setup, where the camera equipped on the UAV such as [19, 14], and 3) Infrared image UAV datasets, including single-frame datasets [6, 4, 5], and video-based Anti-UAV datasets [8, 9, 29, 30] which has been featured in four major challenge events.
@@ 2. $2$
    Provide the experiments validating PIC’s efficiency in terms of IoU and runtime, along with a comprehensive benchmark of UAVDB using SOTA YOLO-series detectors, including YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22].
@@ 3.1.2 Iterative Expansion
At each step $t$, the bounding box expands outward by a fixed size $\delta$ in all directions:
$w_{t+1}=w_{t}+\delta,\quad h_{t+1}=h_{t}+\delta,\quad t=0,1,\ldots$
@@ 3.1.3 Patch Intensity Calculation
The mean pixel intensity at each step inside the bounding box is computed as:
@@ 4.1 Annotation Accuracy and Runtime Efficiency
Here, human-labeled bounding boxes serve as ground truth annotations. For fixed-size and thresholding [1] approaches, we use a 50$\times$50 region and set the threshold to 150 based on empirical tuning for optimal performance. GrabCut [20] and SAM [13] using the OpenCV package and the ViT-B pre-trained model, respectively. As shown in Tab. 3, the PIC approach achieves the highest IoU while maintaining a minimal runtime of 0.007 seconds, comparable to the fixed-size method. This demonstrates that the computational time of the PIC process is negligible compared to image reading and output. In contrast, human labeling takes an average of 19 seconds per annotation, making it impractical for large datasets with tiny objects. Moreover, despite its advanced segmentation capabilities, SAM struggles with UAV-specific challenges, resulting in the lowest IoU. This illustrates that SAM cannot generalize effectively without retraining on a specific dataset. These results highlight the effectiveness of PIC in providing both accurate and computationally efficient UAV bounding box extraction, making it ideal for large-scale and real-time applications.
@@ 4.2 Benchmark on UAVDB
We examine YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22] to benchmark the proposed UAVDB. The experiments were conducted on a high-performance computing (HPC) system [16], utilizing an NVIDIA A100 GPU with 80 GB of memory. All models were trained with an image size of 640, a batch size of 32, 100 epochs, and eight workers. Mosaic augmentation was applied throughout training, excluding the final 10 epochs. Additionally, we fine-tuned the models using officially released pre-trained weights. Tab. 4 summarizes the training time, inference time, model parameters, FLOPs, and average precision (AP) for both validation and test sets. Further, the performance of each model on the validation set across epochs is illustrated in Fig. 4.
@@
Model Training Time Inference Time #Param. (M) FLOPs (G) $\text{AP}^{val}_{50}$ $\text{AP}^{val}_{50-95}$ $\text{AP}^{test}_{50}$ $\text{AP}^{test}_{50-95}$ (hours:mins:sec) (per image, ms) YOLOv8n 01:40:31 0.9 2.685 6.8 0.829 0.522 0.789 0.450 YOLOv8s 01:55:05 1.2 9.828 23.3 0.814 0.545 0.796 0.450 YOLOv8m 02:43:08 1.8 23.203 67.4 0.809 0.538 0.827 0.526 YOLOv8l 03:54:44 2.6 39.434 145.2 0.830 0.563 0.836 0.544 YOLOv8x 04:33:08 3.5 61.597 226.7 0.820 0.554 0.728 0.448 YOLOv9t 02:53:11 2.5 2.617 10.7 0.839 0.501 0.848 0.508 YOLOv9s 03:05:02 2.6 9.598 38.7 0.819 0.517 0.834 0.484 YOLOv9m 05:08:28 4.1 32.553 130.7 0.840 0.507 0.858 0.522 YOLOv9c 06:17:08 5.3 50.698 236.6 0.851 0.544 0.851 0.504 YOLOv9e 08:00:05 6.6 68.548 240.7 0.755 0.414 0.768 0.383 YOLOv10n 02:05:39 0.7 2.695 8.2 0.764 0.492 0.731 0.417 YOLOv10s 02:23:03 1.2 8.036 24.4 0.817 0.530 0.823 0.516 YOLOv10m 03:06:59 1.8 16.452 63.4 0.798 0.531 0.821 0.536 YOLOv10b 03:29:18 2.1 20.413 97.9 0.801 0.517 0.760 0.467 YOLOv10l 04:04:22 2.5 25.718 126.3 0.774 0.502 0.842 0.517 YOLOv10x 05:14:07 3.5 31.586 169.8 0.771 0.507 0.693 0.431 YOLO11n 01:50:00 0.9 2.582 6.3 0.847 0.527 0.856 0.539 YOLO11s 02:07:01 1.2 9.413 21.3 0.826 0.553 0.885 0.578 YOLO11m 03:07:40 1.9 20.031 67.6 0.827 0.588 0.843 0.578 YOLO11l 04:09:45 2.4 25.280 86.6 0.810 0.555 0.798 0.517 YOLO11x 05:20:38 3.6 56.828 194.4 0.812 0.560 0.782 0.534 YOLOv12n 02:15:38 1.8 2.557 6.3 0.857 0.544 0.848 0.531 YOLOv12s 02:44:29 2.0 9.231 21.2 0.869 0.566 0.882 0.565 YOLOv12m 03:34:36 2.6 20.106 67.1 0.866 0.567 0.886 0.584 YOLOv12l 05:10:15 3.1 26.340 88.5 0.870 0.584 0.875 0.590 YOLOv12x 06:35:47 3.9 59.045 198.5 0.879 0.576 0.896 0.569
Table 4: Performance of YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22] models trained on UAVDB.
@@
Fig. 5 presents YOLO11s, the model achieves the best balance between precision and speed, predictions on Dataset 5, where scenarios were absent from the training data, demonstrating its ability to handle unseen situations. The detection results closely match the UAV sizes, validating the high fidelity of the bounding box annotations in UAVDB. Incorporating these high-quality predicted bounding boxes into the training set can further enhance the model’s capability to detect UAVs.
Figure 4: Performance of YOLOv8 [11], YOLOv9 [24], YOLOv10 [23], YOLO11 [10], and YOLOv12 [22] on validation set.
@@ 4.3 Discussion
The proposed PIC generates bounding box annotations with the highest IoU while being approximately 2700$\times$ faster than human labeling. Despite this, the UAVDB remains adequate for training detectors, as shown in Fig. 5. Although the PIC method performs well on current datasets, we recognize that low-altitude UAV flights, with cluttered and rapidly changing backgrounds, may pose challenges. In such dynamic environments, local intensity variations could affect the pixel intensity metric for expanding the bounding box. However, the adaptive nature of PIC, focusing on local intensity changes, allows it to handle moderate variations in background texture. Further improvements for highly dynamic scenarios could include incorporating multi-frame temporal information or background subtraction to enhance robustness and maintain consistent performance.
@@ 5 Conclusion
In this paper, we introduced UAVDB, a dataset with precise bounding box annotations facilitated by our proposed PIC technique. PIC offers an intuitive, efficient, and innovative approach to spatial annotation, eliminating the need for manual labeling. UAVDB addresses critical limitations in existing UAV datasets, such as imprecise annotations and limited environmental diversity, significantly improving the applicability of detection algorithms in real-world scenarios. Through IoU and runtime evaluations for PIC and benchmarking with YOLO-series detectors on UAVDB, we demonstrated the versatility of both UAVDB and PIC approaches under varied conditions. These results establish UAVDB as a valuable resource for advancing UAV detection technologies. Looking ahead, PIC’s adaptability opens promising directions for future research. Its lightweight design could be further optimized by incorporating multi-frame temporal information or background subtraction to improve robustness in dynamic environments. Moreover, its flexibility allows fine-tuning for specific domains, ensuring scalability across various UAV detection applications. As UAV detection technology evolves, UAVDB and PIC provide a solid foundation for advancing real-time, large-scale UAV detection in diverse environments.